# M3RS: Multi-robot, Multi-objective, and Multi-mode Routing and Scheduling

### v2

@@ III-A1 Task Assignment Constraint
[…]
Here, $i=0$ is excluded ($\mathcal{T}\backslash 0$) as it corresponds to the home depot. This is done to maintain feasibility as all agents must leave the home depot. There might be missions where some tasks must be completed. For such cases, the inequality can be replaced by equality for tasks that must be completed.
@@ III-A2 Feasibility Constraints
[…]
@@ III-A3 Energy Consumption Constraint
$\text{q}_{\text{idle}}^{k}+\sum_{i\in\mathcal{T}}(\sum_{m\in\mathcal{M}_{i}}q_{i,m}\cdot y^{k}_{i,m}+\sum_{j\in\mathcal{T}}q_{i\scriptsize{\xrightarrow{}}j}\cdot x^{k}_{i{\xrightarrow{}}j})\,\leq\,Q\,\,\,\forall k\in\mathcal{A}.$ (9)
@@ III-A4 Time Constraints
Given by Eq. (10)-(12), the time constraints ensure the timing of the tasks are consistent:
$s^{k}_{i}+\sum_{m\in\mathcal{M}_{i}}\sigma_{i,m}\cdot y^{k}_{i,m}+t_{i,j}-s^{k}_{j}\leq 0\,\,\forall i\in\mathcal{T}$ (10)
$\forall j\in\mathcal{T}\,\,\forall k\in\mathcal{A},$
@@
The constraint in Eq. (10) ensures that the arrival time for a subsequent task is always later than the previous task. This constraint also enables sub-tour elimination [24], where sub-tours are infeasible solutions that represent disjoint sequences of completing tasks for the same agent. The constraint in Eq. (11) ensures that the agent arrives and completes the task in the selected mode within the task’s prescribed time window. Finally, the constraint in Eq. (12) ensures that all agents return to the depot within the specified time horizon $T_{H}$.
@@ III-B A Solution for M3RS: Clustering-based Column Generation
M3RS extends traditional routing and scheduling problems by adding the decision of mode selection, which impacts task quality. Routing and scheduling problems are known to be $\mathcal{NP}-$hard [25]. Consequently, larger problem instances with many tasks and agents lead to exponential growth in the solution space and significantly burden computational resources [26]. We use CG [27, 28], to address the computational challenges associated with the large decision variables in the proposed formulation. The strength of CG lies in its ability to break down a complex problem into manageable sub-problems while iteratively refining the solution.
The CG algorithm decomposes the original optimization problem into a master problem and a sub-problem. The master problem identifies optimal columns (or paths/solutions) from a pool of feasible candidate solutions, $\Omega$. Since $\Omega$ is very large, typically, a restricted master problem is solved on a subset $\Omega^{\prime}\subset\Omega$. The sub-problem uses the dual solution from the master problem to find new favorable columns, which are added to $\Omega^{\prime}$ to improve the objective value of the master problem. The master problem and sub-problem are solved iteratively until there are no new columns with positive reduced cost. We refer to our approach as CCG (clustering-based CG) due to the use of clustering-based heuristics for solving the sub-problem. The rest of the section presents the building blocks of our CCG algorithm.
@@ III-B1 Restricted Master Problem
This problem aims to identify the best set of solutions from the existing solution pool, resembling a set covering problem. Therefore, to solve the restricted master problem, we adopt a set covering problem formulation [27]. A solution for the set covering problem selects a subset of sets (in our case, task sequence and modes for each agent) from a collection (the solution pool $\Omega^{\prime}$) such that most of the elements (tasks) are covered. This must be done while adhering to constraints, such as each task being assigned to only one agent, and maximizing the given objective. The set covering formulation-based restricted master problem finds optimal solutions from a subset of feasible candidate solutions $\Omega^{\prime}$. Each candidate in $\Omega^{\prime}$ is a feasible solution that specifies a sequence of addressing different tasks in the associated modes and arrival times. The pool $\Omega^{\prime}$ can be initialized by some trivial solutions and can be updated iteratively with better solutions obtained by solving the sub-problem. The set covering formulation for our restricted master problem is given by:
@@
$\begin{split}\sum_{k\in\Omega^{\prime}}a^{k}_{i}\cdot\theta^{k}\leq 1\,\,\forall i\in\mathcal{T}/\{0\},\\ \end{split}$ (18)
@@
Here, $\theta^{k}$ is a binary solution selection variable, where $\theta^{k}=1$ if solution $k\in\Omega^{\prime}$ is selected, and $\theta^{k}=0$ otherwise. The objective function in Eq. (17) selects the candidates from $\Omega^{\prime}$ that maximizes the original objective function Eq. (23). The constraint in Eq. (18) ensures that each task in $\mathcal{T}$ is assigned at most once. Here, $a^{k}_{i}$ is 1 if $i^{th}$ task is covered in $k^{th}$ solution. The constraint in Eq. (19) ensures that mode $m$ for task $i$ is used atmost once with $b^{k}_{i,m}$ being 1 if $i^{th}$ task in $k^{th}$ solution is done in mode $m$. The main purpose of Eq. (19) is to identify the modes to explore via dual values. Finally, Eq. (20) ensures the number of candidates selected from $\Omega^{\prime}$ does not exceed the number of available agents. We apply a linear relaxation to the integer program defined in Eq. (17)-(21) by relaxing the constraint in Eq. (21) and allowing $\theta^{k}\in[0,1]$. This relaxation enables faster computation, as algorithms like the simplex method [29] can efficiently solve large-scale linear programs in polynomial time. This is especially useful with a large number of columns in $\Omega^{\prime}$.
III-B2 Clustering-based Sub-problem
@@
In CG, the sub-problem is solved iteratively to populate $\Omega^{\prime}$. Solving the sub-problem efficiently with high-quality candidate solutions enhances the CG’s exploration of the solution space and accelerates convergence to better solutions. However, as the number of tasks and agents increases, the expanded solution space slows down the search for promising candidates. To address this, we introduce a clustering method inspired by cluster-first, route-second heuristics commonly used in routing and scheduling [30]. Our sub-problem solves the MIP problem in Eq. (22) for a single agent at a time, on a restricted set of tasks $\mathcal{T}^{\prime}$. All tasks in $\mathcal{T}$ are clustered based on time windows. We use a K-Means clustering [31] approach to create $n_{c}$ clusters, with $n_{c}$ being a user-defined parameter. Then $\mathcal{T^{\prime}}$ is generated by uniformly sampling each cluster. These modifications allow CCG to converge to higher-quality solutions faster than the baseline, as demonstrated in Section IV.
@@
The sub-problem creates a solution for a single agent setup with a restricted task pool $\mathcal{T}^{\prime}$. The formulation for the sub-problem is:
@@
$\begin{split}\underset{x_{i\scriptsize{\xrightarrow{}}j},y_{i,m}}{\max}\,\sum_{r=0}^{R}\sum_{i\in\mathcal{T}^{\prime}}\sum_{j\in\mathcal{T}^{\prime}}\sum_{m\in\mathcal{M}_{i}}\lambda_{r}\cdot f_{r}(x_{i\scriptsize{\xrightarrow{}}j},y_{i,m},s_{i})-\\ \sum_{i\in\mathcal{T}^{\prime}/\{0\}}(\gamma_{i}\cdot\sum_{j\in\mathcal{T}^{\prime}}x_{i\scriptsize{\xrightarrow{}}j})-\sum_{i\in\mathcal{T}^{\prime}/\{0\}}\sum_{m\in\mathcal{M}_{i}}(\beta_{i}\cdot y_{i,m}),\end{split}$ (22)
@@
subject to constraints Eq. (3)-(16) and the number of agents is restricted to one. The objective function of the sub-problem, Eq. (22), is a modified variant of Eq. (2), tailored for a single-agent setup and augmented with dual variables from the restricted master problem. Here, $\gamma_{i}$ and $\beta_{i,m}$ are non-negative dual of the restricted master problem corresponding to constraints Eq. (18) and Eq. (19) respectively. These dual variables are used to identify promising new candidate solutions. It should be noted that sub-problem The solution of the sub-problem is added to $\Omega^{\prime}$.
Algorithm 1 The clustering-based column generation algorithm
[…]
@@
15:        $\mathcal{S}\leftarrow$ Solve the sub-problem using Eq. (22)
@@
26:     Solve Eq (17)-(20) subject to $\theta^{k}=\{0,1\}$
@@ III-B3 Complete Algorithm
The CCG algorithm is detailed in Alg. 1. In line 1, the algorithm receives problem parameters, user preferences, and key algorithm settings, including the number of clusters ($n_{c}$), the size of the restricted task set ($|\mathcal{T^{\prime}}|$), maximum iterations, and the time limit. In line 2, initial task clusters $\mathcal{T}_{c}^{init}$ are formed by dividing $\mathcal{T}$ into $n_{c}$ clusters based on time windows using K-Means. In line 3, the candidate solution pool $\Omega^{\prime}$ is initialized, with initial solutions generated by randomly selecting tasks and modes while ensuring constraints in Eq. (3)-(16) are satisfied.
From line 5, the restricted master problem and sub-problem are solved iteratively to generate new solutions. The restricted master problem is updated in each iteration based on $\Omega^{\prime}$ (line 6), and the sub-problem’s objective is guided by the duals obtained from the restricted master. A list $\mathcal{T}_{v}$ is initialized to track visited tasks (line 7). The sub-problem is solved iteratively for each agent in $\mathcal{A}$ (lines 8-20).
In the first iteration, task clusters $\mathcal{T}_{c}$ are initialized to $\mathcal{T}_{c}^{init}$ (line 10). For subsequent iterations, clusters are recreated from the remaining tasks (line 12). The restricted task set $\mathcal{T^{\prime}}$ is formed by uniformly sampling clusters from $\mathcal{T}_{c}$ (line 14), after which the sub-problem is solved (line 15). Feasible solutions with positive reduced costs are added to $\Omega^{\prime}$ (line 17), and $\mathcal{T}_{v}$ is updated with tasks corresponding to the best solutions in $\mathcal{S}$ (line 18). This process continues until the stopping criteria—maximum iterations, time limit, or lack of solutions with positive reduced costs—are met (lines LABEL:lst:ln21-23). Upon termination, the restricted master problem is solved to optimality with integrality constraints to obtain the final solution (lines 25-28). A time limit is applied to both the restricted master and sub-problems to maintain time efficiency throughout the process.

### v3

@@ III-A1 Task Assignment Constraint
$\sum_{k\in\mathcal{A}}\sum_{j\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}j}\,\leq\,1\,\,\,\forall i\in\mathcal{T}\backslash 0.$ (2)
Here, $i=0$ is excluded ($\mathcal{T}\backslash 0$) as it corresponds to the home depot. This is done to maintain feasibility, as all agents must leave the home depot. There might be missions where some tasks must be completed. For such cases, where the tasks must be completed, the inequality can be replaced by equality.
@@ III-A2 Feasibility Constraints
$x^{k}_{i{\xrightarrow{}}i}\,=\,0\,\,\forall k\in\mathcal{A}\,\,\forall i\in\mathcal{T},$ (3)
$\sum_{i\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}0}\,=\,1\,\,\,\forall k\in\mathcal{A},$ (4)
$\sum_{j\in\mathcal{T}}x^{k}_{0\scriptsize{\xrightarrow{}}j}\,=\,1\,\,\,\forall k\in\mathcal{A},$ (5)
$\sum_{i\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}h}-\sum_{j\in\mathcal{T}}x^{k}_{h\scriptsize{\xrightarrow{}}j}\,=\,0\,\,\,\forall h\in\mathcal{T},\,\,\forall k\in\mathcal{A},$ (6)
$\sum_{j\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}j}-\sum_{m\in\mathcal{M}_{i}}y^{k}_{i,m}\,=\,0\,\,\forall(k,i)\in\mathcal{A}\cup\mathcal{T}\backslash 0.$ (7)
@@ III-A3 Energy Consumption Constraint
The energy capacity for each agent is given by $Q$ Ah (Ampere-hours). The energy consumption for each agent is divided into performing tasks, traveling, and idling (Zhang et al.,, 2023). A fixed-rate battery model is used to calculate the energy consumption, where energy consumed is proportional to idling time, task execution time, and traveling time, respectively (Zhang et al.,, 2023). Different aspects of energy expenditure can be calculated using:
$\begin{split}q_{i\scriptsize{\xrightarrow{}}j}=q_{\text{travel}}\cdot t_{i,j},\penalty 10000\ \penalty 10000\ \text{\small(i.e. traveling energy)}\\ q_{i,m}=q_{\text{tc}}\cdot\sigma_{i,m},\penalty 10000\ \penalty 10000\ \text{\small(i.e. servicing energy)}\\ q_{\text{idle}}^{k}=q_{\text{idling}}\cdot(M_{k}-W_{k}).\penalty 10000\ \penalty 10000\ \text{\small(i.e. idling energy)}\end{split}$ (8)
Here, $q_{\text{travel}}$ is the current consumption for an agent traveling at a constant speed, $q_{\text{tc}}$ is the current consumption while performing a task, and $q_{\text{idling}}$ is the current consumption for idling. $M_{k}$ is the total time spent by agent $k$ on the mission, and $W_{k}$ is the total time spent by agent $k$ traveling and executing tasks. The net energy consumed by agent $k$ in a mission is given by the LHS of Eq. (9). The relevant notation for the formulation is summarized in Table I.
@@
$q_{\text{idle}}^{k}+\sum_{i\in\mathcal{T}}(\sum_{m\in\mathcal{M}_{i}}q_{i,m}\cdot y^{k}_{i,m}+\sum_{j\in\mathcal{T}}q_{i\scriptsize{\xrightarrow{}}j}\cdot x^{k}_{i{\xrightarrow{}}j})\,\leq\,Q\,\,\,\forall k\in\mathcal{A}.$ (9)
@@ III-A4 Time Constraints
Given by Eq. (10)-(12), the time constraints ensure the timing of the tasks is consistent:
$s^{k}_{i}+\sum_{m\in\mathcal{M}_{i}}\sigma_{i,m}\cdot y^{k}_{i,m}+t_{i,j}-s^{k}_{j}\leq 0\,\,\forall i\in\mathcal{T}\\ \forall j\in\mathcal{T}\,\,\forall k\in\mathcal{A},$ (10)
@@
The constraint in Eq. (10) ensures that the arrival time for a subsequent task is always later than the previous task. This constraint also enables sub-tour elimination (Kara et al.,, 2004), where sub-tours are infeasible solutions that represent disjoint sequences of completing tasks for the same agent. The constraint in Eq. (11) ensures that the agent arrives and completes the task in the selected mode within the task’s prescribed time window. Finally, the constraint in Eq. (12) ensures that all agents return to the depot within the specified time horizon $T_{H}$.
@@ III-B A Solution for M3RS: Clustering-based Column Generation
M3RS extends traditional routing and scheduling problems by incorporating the decision of mode selection, which affects task quality. Routing and scheduling problems are known to be $\mathcal{NP}$-hard (Kumar and Panneerselvam,, 2012). Consequently, larger problem instances with many tasks and agents lead to exponential growth in the solution space and significantly burden computational resources (Uchoa et al.,, 2017). We use Column Generation (CG) (Desrosiers and Lübbecke,, 2005; Lübbecke and Desrosiers,, 2005) to address the computational challenges associated with the large decision variables in the proposed formulation. The strength of CG lies in its ability to break down a complex problem into manageable subproblems while iteratively refining the solution.
The CG algorithm decomposes a large-scale optimization problem into two components: a master problem and a subproblem (see Fig. 2 (a)). The master problem, typically formulated as a linear program (Vanderbeck,, 2005), selects the best combination of columns (also referred to as paths or candidate solutions) from a feasible set $\Omega$. Since $\Omega$ is generally too large to consider in full, a restricted master problem is solved over a smaller subset $\Omega^{\prime}\subset\Omega$. Solving this restricted problem yields a tentative solution along with dual variables. In a linear program, each constraint in the primal problem corresponds to a variable in the dual problem, known as a dual variable (Bachem et al.,, 1992). Intuitively, a dual variable represents the rate at which the optimal value of the primal objective would improve if the associated constraint were relaxed by one unit. The subproblem uses these dual variables to identify new columns that could improve the current solution. Guided by the duals, the subproblem searches $\Omega$ for high-value candidates missing from $\Omega^{\prime}$. These candidates are added to $\Omega^{\prime}$, and the master problem is re-solved. For example, the dual variables associated with the restricted master problem constraints in Eqs. (18)–(19) are used in the objective function of the subproblem defined in Eq. (22). This iterative process continues until no further columns can be found that would improve the objective value of the master problem.
The overall performance of the CG algorithm heavily depends on how effectively the subproblem is solved, particularly as the number of agents and tasks increases. In such cases, identifying promising solution candidates to expand the restricted set $\Omega^{\prime}$ becomes computationally challenging. To address this limitation, we propose a clustering-based heuristic method, coined Clustering-based Column Generation (CCG), to solve the subproblem method efficiently. CCG tackles the subproblem one agent at a time, focusing on a small, representative subset of tasks, denoted by $\mathcal{T}^{\prime}$ (see Fig. 2 (b)). This subset is constructed by first clustering the original task set $\mathcal{T}$ based on task time-windows, and then uniformly sampling from these clusters to form $\mathcal{T}^{\prime}$. This targeted reduction in task space allows for more scalable and efficient subproblem solving. The detailed methodology behind CCG is presented in the following section.
Master ProblemDual ValuesSubproblemSubset SolutionPool ($\Omega^{\prime}\subset\Omega$)Solution Pool ($\Omega$)
(a) Block diagram of Column Generation (CG)
Master ProblemSubset SolutionPool ($\Omega^{\prime}\subset\Omega$)Solution Pool ($\Omega$)Dual ValuesAvailable TasksCreate ClustersSingle-agentMIPSampled task subset($\mathcal{T^{\prime}}$)SubproblemTask Set ($\mathcal{T}$)
(b) Block diagram of the proposed Clustering Column Generation (CCG)
Figure 2: (a) Column Generation (CG) decomposes an optimization problem into a master problem and a subproblem. The master problem selects the best combination of solutions from a solution pool ($\Omega$), while the subproblem uses dual values from the master to generate promising new candidates. Since $\Omega$ can be very large the CG operates on a smaller subset of solutions, i.e. $\Omega\subset\Omega^{\prime}$. This process iterates until no improving solutions are found or a termination condition is met. (b) We propose Clustering-based Column Generation (CCG), which incorporates a clustering heuristic to efficiently solve the subproblem. To reduce computational cost, tasks in the orignal task set $\mathcal{T}$, are first grouped into clusters using $k$-means based on time windows. For each agent, a MIP with a time limit is solved on a reduced subset of tasks $\mathcal{T^{\prime}}$, obtained by uniformly sampling from these clusters. The resulting solutions are added to the subset solution pool ($\Omega^{\prime}$) for use in the master problem.
@@ III-B1 Restricted Master Problem
This problem aims to identify the best set of solutions from the existing solution pool, resembling a set covering problem (Caprara et al.,, 2000). Therefore, to solve the restricted master problem, we adopt a set covering problem formulation (Desrosiers and Lübbecke,, 2005). A solution for the set covering problem selects a subset of sets (in our case, task sequences and modes for each agent) from a collection (the solution pool $\Omega^{\prime}$) such that most of the elements (tasks) are covered. This must be done while adhering to constraints, such as each task being assigned to only one agent and maximizing the given objective. The set covering formulation-based restricted master problem finds optimal solutions from a subset of feasible candidate solutions $\Omega^{\prime}$. Each candidate in $\Omega^{\prime}$ is a feasible solution that specifies a sequence of addressing different tasks in the associated modes and arrival times. The pool $\Omega^{\prime}$ can be initialized by some trivial solutions and can be updated iteratively with better solutions obtained by solving the subproblem. The set covering formulation for our restricted master problem is given by:
@@
[…]
@@
Here, $\theta^{k}$ is a binary solution selection variable, where $\theta^{k}=1$ if solution $k\in\Omega^{\prime}$ is selected, and $\theta^{k}=0$ otherwise. The objective function in Eq. (17) selects the candidates from $\Omega^{\prime}$ that maximize the original objective function Eq. (23). The constraint in Eq. (18) ensures that each task in $\mathcal{T}$ is assigned at most once. Here, $a^{k}_{i}$ is 1 if $i^{th}$ task is covered in $k^{th}$ solution. The constraint in Eq. (19) ensures that mode $m$ for task $i$ is used at most once with $b^{k}_{i,m}$ being 1 if $i^{th}$ task in $k^{th}$ solution is done in mode $m$. The main purpose of Eq. (19) is to identify the modes to explore via dual values. Finally, Eq. (20) ensures the number of candidates selected from $\Omega^{\prime}$ does not exceed the number of available agents. Typically, a linear relaxation of the original restricted master problem is solved in column generation (Vanderbeck,, 2005). This is because the dual of a mixed integer programming (MIP) problem is not well defined, which poses a challenge since dual variables are essential for solving the subproblem. In contrast, dual variables are well defined for linear programs (Guzelsoy and Ralphs,, 2007). The linear relaxation of the integer program defined in Eqs. (17)–(21) is obtained by relaxing the integrality constraint in Eq. (21) and allowing $\theta^{k}\in[0,1]$. This relaxation also enables faster computation, as algorithms such as the simplex method (Nabli,, 2009) can efficiently solve large-scale linear programs. This is particularly advantageous when the restricted column set $\Omega^{\prime}$ contains a large number of variables.
III-B2 Clustering-based subproblem
@@
In CG, the subproblem is solved iteratively to populate $\Omega^{\prime}$. Solving the subproblem efficiently with high-quality candidate solutions enhances the CG’s exploration of the solution space and accelerates convergence to better solutions. However, as the number of tasks and agents increases, the expanded solution space slows down the search for promising candidates. To address this, we introduce a clustering method inspired by cluster-first, route-second heuristics commonly used in routing and scheduling (Liu et al.,, 2023; Elango et al.,, 2011). Our subproblem solves the MIP problem in Eq. (22) for a single agent at a time, on a restricted set of tasks $\mathcal{T}^{\prime}$. All tasks in $\mathcal{T}$ are clustered based on time windows. We use a $k$-means clustering (Bishop,, 2006) approach to create $n_{c}$ clusters, with $n_{c}$ being a user-defined parameter. Then $\mathcal{T^{\prime}}$ is generated by uniformly sampling each cluster. These modifications allow CCG to converge to higher-quality solutions faster than the baseline, as demonstrated in Section IV.
@@
The subproblem creates a solution for a single-agent setup with a restricted task pool $\mathcal{T}^{\prime}$. The subproblem formulation is:
@@
$\begin{split}\underset{x_{i\scriptsize{\xrightarrow{}}j},y_{i,m}}{\max}\,\sum_{r=0}^{R}\sum_{i\in\mathcal{T}^{\prime}}\sum_{j\in\mathcal{T}^{\prime}}\sum_{m\in\mathcal{M}_{i}}\lambda_{r}\cdot f_{r}(x_{i\scriptsize{\xrightarrow{}}j},y_{i,m},s_{i})-\sum_{i\in\mathcal{T}^{\prime}/\{0\}}(\gamma_{i}\cdot\sum_{j\in\mathcal{T}^{\prime}}x_{i\scriptsize{\xrightarrow{}}j})-\sum_{i\in\mathcal{T}^{\prime}/\{0\}}\sum_{m\in\mathcal{M}_{i}}(\beta_{i}\cdot y_{i,m}),\end{split}$ (22)
@@
subject to constraints Eq. (2)-(16) and the number of agents is restricted to one. The objective function of the subproblem, Eq. (22), is a modified variant of Eq. (1), tailored for a single-agent setup and augmented with dual variables from the restricted master problem. Here, $\gamma_{i}$ and $\beta_{i,m}$ are non-negative duals of the restricted master problem corresponding to constraints Eq. (18) and Eq. (19) respectively. These dual variables are used to identify promising new candidate solutions. Dual variables can be directly obtained from the LP relaxation solution provided by the optimizer (Laborie et al.,, 2018). It should be noted that the solution of the subproblem is added to $\Omega^{\prime}$.
Algorithm 1 Clustering-based Column Generation (CCG)
1:  Input: Problem data ($\mathcal{A},\mathcal{T},\mathcal{M},T_{H},Q$ and associated data for nodes), user preference $\lambda$, and algorithm parameters ($n_{c},|\mathcal{T^{\prime}}|,\text{max-iterations},\text{time-limit}$).
@@
15:    $\mathcal{S}\leftarrow$ Solve the subproblem using Eq. (22)
@@
26:   Solve Eq. (17)-(20) subject to $\theta^{k}=\{0,1\}$
@@ III-B3 Complete Algorithm
The CCG algorithm is detailed in Alg. 1. In line 1, the algorithm receives problem parameters, user preferences, and key algorithm settings, including the number of clusters ($n_{c}$), the size of the restricted task set ($|\mathcal{T^{\prime}}|$), maximum iterations, and the time limit. In line 2, initial task clusters $\mathcal{T}_{c}^{init}$ are formed by dividing $\mathcal{T}$ into $n_{c}$ clusters based on time windows using $k$-means. In line 3, the candidate solution pool $\Omega^{\prime}$ is initialized, with initial solutions generated by randomly selecting tasks and modes while ensuring constraints in Eq. (2)-(16) are satisfied. From line 5, the restricted master problem and subproblem are solved iteratively to generate new solutions. The restricted master problem is updated in each iteration based on $\Omega^{\prime}$ (line 6), and the subproblem’s objective is guided by the value of dual variables obtained from the restricted master. A list $\mathcal{T}_{v}$ is initialized to track visited tasks (line 7). The subproblem is solved iteratively for each agent in $\mathcal{A}$ (lines 8-20). In the first iteration, task clusters $\mathcal{T}_{c}$ are initialized to $\mathcal{T}_{c}^{init}$ (line 10). For subsequent iterations, clusters are recreated from the remaining tasks (line 12). The restricted task set $\mathcal{T^{\prime}}$ is formed by uniformly sampling clusters from $\mathcal{T}_{c}$ (line 14), after which the subproblem is solved (line 15). Feasible solutions with positive reduced costs are added to $\Omega^{\prime}$ (line 17), and $\mathcal{T}_{v}$ is updated with tasks corresponding to the best solutions in $\mathcal{S}$ (line 18). This process continues until the stopping criteria—maximum iterations, time limit, or lack of solutions with positive reduced costs—are met (lines 21-23). Upon termination, the restricted master problem is solved to optimality with integrality constraints to obtain the final solution (lines 25-28). A time limit is applied to the restricted master and subproblems to maintain time efficiency during the process.
The worst-case time complexity of CCG is determined by its components. The restricted master problem is a linear program, which, when solved using the simplex method, has a worst-case exponential complexity of $O(2^{m})$ (Voulgaropoulou et al.,, 2019), where $m$ is the number of constraints. However, it typically runs in polynomial time in practice (Voulgaropoulou et al.,, 2019). The subproblem, solved as an MIP using a branch-and-bound solver, also possesses exponential complexity (Huang et al.,, 2021). Additionally, the $k$-means clustering step has a worst-case complexity of $O(n^{d})$, where $n$ is the number of nodes and $d$ is the data dimension (Arthur and Vassilvitskii,, 2006). However, in practical implementations, $k$-means typically converges in $O(nkT)$ time (Arthur and Vassilvitskii,, 2006), where $k$ is the number of clusters and $T$ is the number of iterations. With a fixed iteration limit, its runtime becomes effectively $O(1)$ per instance. Since both the master and subproblem have fixed time limits, their complexity per iteration is effectively $O(1)$. With an efficient implementation of $k$-means that also runs in bounded time, the overall time complexity is reduced to $O(I)$, where $I$ is the number of iterations until the convergence or time-out of the CCG. The space complexity depends on the size of the solution pool, leading to $O(|\Omega^{\prime}|)$.